OpenStudy (camerondoherty):

A system of equations is shown below: x + y = 3 2x – y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (camerondoherty):

im thinking 3

OpenStudy (dan815):

you can add the 2 equations, or use method of substitution

OpenStudy (dan815):

x + y = 3 2x – y = 6 ------------ 3x +0y = 9 x=9/3=3

OpenStudy (dan815):

right again

OpenStudy (camerondoherty):

Thank You!